Westerville Company reported the following results from last years operations: Sales Variable expenses Contribution margin F4. What is the margin related to this years investment opportunity? Margin %6. What is the ROI related to this years investment opportunity? (Do not round intermediate calculations.) ROL7. If the company pursues the investment opportunity and otherwise performs the same as last year, what margin will it earn t8. If the company pursues the investment opportunity and otherwise performs the same as last year, what turnover will it earn9. If the company pursues the investment opportunity and otherwise performs the same as last year, what ROI will it earn this
Answer #1

Margin = 420000*70%-252000 = 42000

4) Margin = 42000/420000 = 10%

6) ROI = 42000/350000 = 12%

7) Margin = (375000+42000)/1920000 = 21.7%

8) Turnover = 1920000/1600000 = 1.20

9) ROI = 417000/1600000 = 26.1%